# Inverse Virtual Try-On: Generating Multi-Category Product-Style Images from Clothed Individuals

### v1

@@ 5 Conclusion
We have presented TEMU-VTOFF, a novel virtual try-off architecture designed to generate high-quality, diverse in-shop garment images. Our approach is the first to effectively extend VTOFF to multi-category scenarios, addressing existing inherent shortcomings of existing methods with ad-hoc modules such as a dual-DiT architecture equipped with our multimodal hybrid attention to effortlessly integrate garment, person, and text features. Moreover, we have proposed an external garment aligner module paired with a new loss objective for visual detail refinement. Experimental results on standard VTOFF benchmarks demonstrate the robustness and effectiveness of our approach across a range of garment types and evaluation metrics.
[…]
Appendix A Experimental Protocols
@@ A.1 Datasets Details
Dress Code. In our experiments, we adopt the Dress Code dataset [12], the largest publicly available benchmark for image-based virtual try-on. Unlike previous datasets limited to upper-body clothing, Dress Code includes three macro-categories:
  • Upper-body: $15,363$ pairs (e.g., tops, t-shirts, shirts, sweatshirts)
  • Lower-body: $8,951$ pairs (e.g., trousers, skirts, shorts)
  • Dresses: $29,478$ (e.g., full body dresses)
The total number of paired samples is $53,792$, split into $48,392$ training images and $5,400$ test images at a resolution of $1024\times 768$.
VITON-HD. Following previous literature, we also adopt VITON-HD [8], a publicly available dataset widely used in virtual try-on research. It is composed exclusively of upper-body garments and provides high-resolution images at $1024\times 768$ pixels. The dataset contains a total of $27,358$ images, structured into $13,679$ garment-model pairs. These are split into $11,647$ training pairs and $2,032$ test pairs, each comprising a front-view image of a garment and the corresponding image of a model wearing it.
@@ A.2 Implementation Details
We evaluate our method both with distribution-based metrics and per-sample similarity metrics. For the first group, we adopt FID [37] and KID [4] implementations derived from clean-fid PyTorch packagehttps://pypi.org/project/clean-fid/. Concerning the second group, we adopt both SSIM [50] and LPIPS [58] as they are the standard metrics adopted in the field to measure structural and perceptual similarity between a pair of images. We reuse the corresponding Python packages provided by TorchMetricshttps://pypi.org/project/torchmetrics/. Finally, we adopt DISTS [13] as an additional sample-based similarity metric, as it correlates better with human judgment, as shown in previous works [16]. We stick to the corresponding Python package§§§https://pypi.org/project/DISTS-pytorch/ to compute it for our experiments.
A.3 Algorithm
@@
Algorithm 1 TEMU-VTOFF: Virtual Try-Off with Dual-DiT and Garment Alignment
@@
11:$\bm{e}^{v}_{\text{pool}}\leftarrow\text{CLIP}(\bm{x}_{\text{model}})$
@@
13:$\{\bm{K}^{l}_{\text{extractor}},\bm{V}^{l}_{\text{extractor}}\}_{l=1}^{N}\leftarrow F_{E}(\bm{z}^{\prime}_{0},\bm{x}_{\text{model}},t{=}0)$
@@
18:$\hat{\epsilon}_{t}\leftarrow F_{D}(\bm{z}_{t},\bm{e}_{\text{pooled}},\bm{e}_{\text{text}},\{\bm{K}^{l}_{\text{extractor}},\bm{V}^{l}_{\text{extractor}}\},t)$
19: Compute diffusion loss: $\mathcal{L}_{\text{diff}}\leftarrow\|\hat{\epsilon}_{t}-\epsilon_{t}\|^{2}$
@@
26: Combine losses: $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{diff}}+\lambda\cdot\mathcal{L}_{\text{align}}$
@@
Appendix B Caption Extraction Details
We leverage Qwen2.5-VL [1] to generate the caption of a given garment image, following the chat template provided below:
visual_attributes = {
"lower_body": ["Cloth Type", "Waist", "Fit", "Cloth Length"]
category = "upper_body" # The category to which the image belongs
User:[Uncaptioned image] Use only visual attributes that are present in the image. Predict values of the following attributes: {visual_attributes[category]}. It’s forbidden to generate the following visual attributes: colors, background, and textures/patterns. It’s forbidden to generate unspecified predictions. It’s forbidden to generate newline characters. Generate in this way: a <cloth type> with <attributes description>
Appendix C Additional Qualitative Results
@@
We show an extended version of the qualitative results presented in our main paper. We report additional visual results on the competitors in Fig. F on the Dress Code [12] dataset and in Fig. G on the VITON-HD [8] dataset. Furthermore, Fig. H presents additional ablation results to analyze the impact of textual and mask conditioning. Finally, we include in Fig. I the full set of inputs used for generating the target garment, including the model input, the segmentation mask, and the caption.
Figure F: Additional qualitative results of TEMU-VTOFF and competitors on Dress Code [12].
Figure G: Additional qualitative results of TEMU-VTOFF and competitors on VITON-HD [8].
Figure H: Additional qualitative results showing the contribution of each component in TEMU-VTOFF on Dress Code [12] images.
Figure I: Inputs used to generate the target garment with TEMU-VTOFF, using sample images from Dress Code [11]
Appendix D Limitations
@@
Our model inherits some inner problems of foundational models like Stable Diffusion 3 [14]. Even though our method improved the generation of big logos and text, it is limited in its scope concerning fine-grained details like complex texture patterns and small written text. Moreover, it sometimes fails to render the correct number of small objects like buttons. We show a set of failure cases in Fig. J and Fig. K, on sample images from Dress Code [11] and VITON-HD [32] respectively.
Appendix E Broader Impact
Our method addresses the VTOFF task by generating flat, in-shop garment images from photos of dressed individuals. This enables a novel form of data augmentation in the fashion domain, allowing clean garment representations to be synthesized without manual segmentation or dedicated photoshoots. By bridging the gap between worn and catalog-like appearances, our approach can improve scalability for fashion datasets and support downstream applications such as retrieval, recommendation, and virtual try-on.
However, as with any generative technology, there are important ethical and legal considerations. In particular, our model could be used to reconstruct garments originally designed by third parties, potentially raising issues of copyright and intellectual property infringement. We emphasize that our framework is intended for research and responsible use, and any deployment in commercial settings should ensure compliance with applicable copyright laws and respect for designer rights.

### v2

@@ 5 Conclusion
We presented TEMU-VTOFF, a novel architecture that pushes the boundaries of VTOFF for complex, multi-category scenarios. While existing methods often struggle with detail preservation and accurate reconstruction across diverse garment types, our approach is specifically designed to overcome these limitations. We achieve this through a novel dual-DiT framework that leverages multimodal hybrid attention to effectively fuse information from the person, the garment, and textual descriptions. To enhance realism, our proposed garment aligner module refines fine-grained textures and structural details. The effectiveness of our method is validated by state-of-the-art performance on standard VTOFF benchmarks, demonstrating its robustness in generating high-fidelity, catalog-style images.
Acknowledgments. This work was supported by EU Horizon projects ELIAS (No. 101120237) and ELLIOT (No. 101214398), and by the FIS project GUIDANCE (No. FIS2023-03251).
Ethics Statement
Our method addresses the VTOFF task by generating flat, in-shop garment images from photos of dressed individuals. This enables a novel form of data augmentation in the fashion domain, allowing clean garment representations to be synthesized without manual segmentation or dedicated photoshoots. By bridging the gap between worn and catalog-like appearances, our approach can improve scalability for fashion datasets and support downstream applications such as retrieval, recommendation, and virtual try-on. However, as with any generative technology, there are important ethical and legal considerations. In particular, our model could be used to reconstruct garments originally designed by third parties, potentially raising issues of copyright and intellectual property infringement. We emphasize that our framework is intended for research and responsible use, and any deployment in commercial settings should ensure compliance with applicable copyright laws and respect for designer rights.
Reproducibility Statement
This work uses only public datasets and open-source models for its training and evaluations. In the Appendix, we include all the implementation and dataset details to reproduce our results. In addition, we will publicly release the source code and trained models to further support reproducibility.
A Additional Details
@@ A.1 Datasets Details
Dress Code. In our experiments, we adopt the Dress Code dataset (Morelli et al., 2022), the largest publicly available benchmark for image-based virtual try-on. Unlike previous datasets limited to upper-body clothing, Dress Code includes three macro-categories: upper-body clothes with 15,363 pairs (e.g., tops, t-shirts, shirts, sweatshirts), lower-body clothing with 8,951 pairs (e.g., trousers, skirts, shorts), and full-body dresses with 29,478 pairs. The total number of paired samples is $53,792$, split into $48,392$ training images and $5,400$ test images at a resolution of $1024\times 768$.
VITON-HD. Following previous literature, we also adopt VITON-HD (Choi et al., 2021), a publicly available dataset widely used in virtual try-on research. It is composed exclusively of upper-body garments and provides high-resolution images at $1024\times 768$ pixels. The dataset contains a total of $27,358$ images, structured into $13,679$ garment-model pairs. These are split into $11,647$ training pairs and $2,032$ test pairs, each comprising a front-view image of a garment and the corresponding image of a model wearing it.
@@ A.2 Implementation Details
For both the feature extractor and the diffusion backbone, we adopt Stable Diffusion 3 medium (Esser et al., 2024). All models are trained on a single node equipped with 4 NVIDIA A100 GPUs (64GB each), using DeepSpeed ZeRO-2 (Rajbhandari et al., 2021) for efficient distributed training. We use a total batch size of 32 and train each model for 30k steps, corresponding to approximately 960k images. Optimization is performed with AdamW (Loshchilov and Hutter, 2019), using a learning rate of $1\times 10^{-4}$, a warmup phase of 3k steps, and a cosine annealing schedule. We train separate models per dataset to account for differences in distribution and garment structure. In all experiments, we set the alignment loss weight $\lambda$ equal to $0.5$.
We evaluate our method both with distribution-based metrics and per-sample similarity metrics. For the first group, we adopt FID (Parmar et al., 2022) and KID (bińkowski2021demystifyingmmdgans) implementations derived from clean-fid PyTorch package [^2] https://pypi.org/project/clean-fid/. Concerning the second group, we adopt both SSIM (Wang et al., 2004), PSNR (Wang et al., 2004), and LPIPS (Zhang et al., 2018) as they are the standard metrics adopted in the field to measure structural and perceptual similarity between a pair of images. We reuse the corresponding Python packages provided by TorchMetrics [^3] https://pypi.org/project/torchmetrics/. Finally, we adopt DISTS (Ding et al., 2020) as an additional sample-based similarity metric, as it correlates better with human judgment, as shown in previous works (Fu et al., 2023). We stick to the corresponding Python package [^4] https://pypi.org/project/DISTS-pytorch/ to compute it for our experiments.
A.3 Caption Extraction Details
We leverage Qwen2.5-VL (Bai et al., 2025) to generate textual descriptions of the garments, which serve as semantic conditioning for our Dual-DiT architecture.
To ensure no ground-truth information leaks into the testing process, we employ two different generation pipelines for training and testing:
  • Training: We generate captions using the ground-truth in-shop garment images. This ensures the model learns precise semantic correlations between visual features and textual attributes during optimization (see Fig. 6).
  • Inference: At test time, the ground-truth garment image is strictly unavailable. Instead, the caption is generated directly from the input person image. Qwen2.5-VL is prompted to analyze the person’s clothing and extract the relevant structural attributes. This ensures our method is fully applicable to “in-the-wild” scenarios where only the person image is known (see Fig. 7).
We define a variable category $\in$ {dress, upper body, lower body}. To avoid leaking color or texture information (which should be handled by the visual feature extractor $F_{E}$) and to focus solely on structural guidance, we utilize a strict prompt template. The prompt explicitly forbids the generation of non-structural attributes:
visual_attributes = {
"lower_body": ["Cloth Type", "Waist", "Fit", "Cloth Length"]
User:Refer to caption Use only visual attributes that are present in the image. Predict values of the following attributes: {visual_attributes[category]}. It’s forbidden to generate the following visual attributes: colors, background, and textures/patterns. It’s forbidden to generate unspecified predictions. It’s forbidden to generate newline characters. Generate in this way: a <cloth type> with <attributes description>.
Figure 6: Caption extraction pipeline (training stage).
User:Refer to caption Use only visual attributes that are present in the image. Predict values of the following attributes: {visual_attributes[category]}. Do it only for the category garment. It’s forbidden to generate the following visual attributes: colors, background, and textures/patterns. It’s forbidden to generate unspecified predictions. It’s forbidden to generate newline characters. Generate in this way: a <cloth type> with <attributes description>.
Qwen Caption: A leather skirt with a fitted waist and a short length.
Figure 7: Caption extraction pipeline (inference stage).
A.4 Algorithm
@@
Algorithm 1 TEMU-VTOFF: Dual-DiT and Garment Alignment for VTOFF
@@
[…]
@@
[…]
@@
[…]
19: Compute diffusion loss: $\mathcal{L}_{\text{DiT}}\leftarrow\|\hat{\epsilon}_{t}-\epsilon_{t}\|^{2}$
@@
26: Combine losses: $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{DiT}}+\lambda\cdot\mathcal{L}_{\text{align}}$
@@
B Additional Quantitative Results and Analyses
In this section, we report additional quantitative results and analyses on the effectiveness of the proposed components and design choices.
Effect of Varying $\lambda$ Parameter. We conducted an ablation study on the Dress Code dataset to assess the effect of the $\lambda$ regularization for the alignment of our main diffusion transformer $F_{D}$ and the DINOv2 features. We report the results in Table 6. As shown, $\lambda=0.5$ is the overall best choice across all metrics.
Effect of Varying the DiT Block $i$ Used for $\mathcal{L}_{align}$. A critical design choice in our garment aligner module is determining which internal block of the DiT ($F_{D}$) should be aligned with the semantic features from DINOv2. To find the optimal depth, we conduct an ablation study on the Dress Code dataset, varying the block index $i\in\{6,8,12,18\}$ within the 24-block SD 3 medium architecture.
The results are reported in the middle section of Table 6. We observe that aligning the 8th block yields the best overall performance. These results highlight a trade-off between structural guidance and generation flexibility, consistent with recent findings in representation alignment (Yu et al., 2025). The early-to-mid layers of the DiT capture the coarse semantic layout and structural essence of the image. Aligning these with DINOv2 ensures the generated garment respects the target structure while leaving subsequent layers free to refine details. As we move to deeper blocks, perceptual quality degrades. While these layers maintain high structural similarity, the distributional metrics worsen. This occurs because the deeper layers of a DiT are increasingly specialized in predicting the high-frequency noise (or flow velocity) required for the diffusion objective. Forcing these “noisy” layers to align with the “clean”, invariant features of DINOv2 introduces an optimization conflict, ultimately smoothing out fine-grained textures and degrading realism.
Analysis of Asynchronous Timestep Conditioning. A critical design choice in our architecture is the use of a fixed timestep $t=0$ for the feature extractor $F_{E}$, while the main denoising DiT $F_{D}$ operates on a noisy latent $z_{t}$ at timestep $t>0$. This raises an important question: could this discrepancy in timesteps lead to a misalignment in the feature space? In this section, we provide the rationale for this design choice, supported by concurrent work and a targeted ablation study.
Our primary motivation is to provide the main generator $F_{D}$ with the cleanest, most semantically rich conditioning signal possible. By extracting features from $F_{E}$ at $t=0$ we ensure the conditioning information is completely free from stochastic noise inherent to the diffusion process. We hypothesize that injecting features from a noisy timestep $t>0$ would introduce an additional, confounding source of noise into the generation process, thereby degrading the quality of the final output. The key to our method is that the MHA module is specifically trained to bridge this temporal gap; it learns to effectively attend to the clean conditioning features to guide the denoising of the noisy latent $z_{t}$.
This design philosophy is strongly supported by recent, concurrent research that analyzes the internal representations of diffusion models:
  • The work on CleanDIFT (Stracke et al., 2025) directly argues that adding noise to images before feature extraction is a performance bottleneck that harms feature quality. Their entire method is built on the same premise as our $F_{E}$: that extracting features from clean images leads to superior performance without needing task-specific timestep tuning.
  • Furthermore, ConceptAttention (Helbling et al., 2025) demonstrates that the internal representations of DiTs are highly interpretable and correspond to semantic concepts, particularly at early timesteps. This validates our choice to use $t=0$ features, as they represent the purest and most semantically meaningful form of this information.
To validate our design choice, we conducted an ablation study comparing our method with the variant where the feature extractor $F_{E}$ and the denoising $F_{D}$ use the same synchronous timestep $t$. The results on Dress Code are presented in Table 6. As shown in the table, our proposed method with asynchronous timesteps significantly outperforms the synchronous variant across the majority of the metrics. This result provides strong empirical evidence for the value of clean conditioning and confirms the effectiveness of our proposed Multimodal Hybrid Attention.
Table 6: Additional ablation study results on the Dress Code dataset.
All Upper-body Lower-body Dresses
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$
Effect of Varying $\lambda$ Parameter
$\lambda=0.0$ (w/o g. aligner) 76.01 12.85 30.84 20.63 5.91 0.78 21.77 11.26 22.26 14.22 17.86 11.86
$\lambda=0.25$ 74.29 12.48 33.68 19.41 6.42 0.89 19.65 10.77 21.86 16.53 16.73 11.38
$\lambda=0.5$ (Ours) 75.95 12.90 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
$\lambda=0.75$ 71.93 11.71 37.03 20.35 7.81 1.41 20.11 11.09 23.85 21.12 17.10 11.69
$\lambda=1.0$ 71.76 11.72 37.21 20.45 7.78 1.39 20.07 11.33 24.11 20.59 17.17 11.79
Effect of Varying the DiT Block $i$ Used for $\mathcal{L}_{align}$
$i=6$ 72.11 11.44 38.00 20.62 8.66 1.76 21.21 12.36 23.25 22.15 17.42 12.49
$i=8$ (Ours) 75.95 12.90 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
$i=12$ 75.30 12.59 32.71 19.13 6.48 0.87 20.41 11.86 20.01 15.40 16.98 11.72
$i=18$ 76.16 12.57 31.66 19.17 6.87 1.16 20.27 12.37 19.86 15.52 17.38 12.28
Effect of Asynchronous Timestep Conditioning
w/ same $t$ in $F_{E}$ and $F_{D}$ 77.70 12.66 32.69 22.41 9.78 2.30 23.98 17.85 21.29 17.83 21.95 17.52
w/ $t=0$ in $F_{E}$ (Ours) 75.95 12.90 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
Generalization with Stronger Vision Encoders. We replace our CLIP vision encoder with a more powerful SigLIP 2 (Tschannen et al., 2025). We adopt the ViT-g 16 model and retrained our architecture with an additional MLP $f_{\mathrm{MLP}}:\mathbb{R}^{1536}\rightarrow\mathbb{R}^{2048}$ to project SigLIP 2 output dimension into the SD3 input space. The results are presented in Table 7. As noted in TryOffDiff (Velioglu et al., 2024), employing a stronger vision encoder improves the final performance. Our experiments further validate this finding. This improvement is due to the better capacity of SigLIP 2 at extracting more fine-grained features. As reported in (Tschannen et al., 2025), this model is trained with a binary contrastive loss that processes each text-image pair separately, thus preventing information corruption from different image-text pairs. Moreover, fine-grained details are enhanced with a self-distillation loss and masked prediction. Finally, this ablation further demonstrates that our core contribution lies in our dual-DiT architecture because this design scheme can be improved with plug-and-play modules, unlike the architectural alternatives that underperform in the same setting.
Table 7: Comparison of CLIP vs. SigLIP 2 as vision encoder for person encoding. $\uparrow$ indicates higher is better, $\downarrow$ lower is better.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TEMU-VTOFF w/ CLIP 75.95 12.90 31.46 18.66 5.74 0.65
TEMU-VTOFF w/ SigLIP 2 76.62 14.33 28.10 18.77 5.08 0.53
C Additional Qualitative Results
@@
We report an extended version of the qualitative results presented in our main paper. Specifically, additional visual comparisons between TEMU-VTOFF and competitors are shown in Fig. 9 and Fig. 10, on sample images from Dress Code (Morelli et al., 2022) and VITON-HD (Choi et al., 2021), respectively. Moreover, Fig. 11 presents additional ablation results to analyze the impact of textual and mask conditioning. Finally, we include in Fig. 12 the full set of inputs used for generating the target garment, including the model input, the segmentation mask, and the textual caption.
D User Study
To complement our quantitative analysis and address the limitations of automated metrics in capturing fine-grained texture details, we conduct a human perceptual study.
Experimental Setup. We recruited 42 distinct participants to evaluate the visual quality of the generated garments. The study followed a pairwise comparison protocol. For each trial, participants were presented with the input person image and two generated garment results: one from TEMU-VTOFF and one from a competitor (randomly selected from MGT (Velioglu et al., 2025) or Any2AnyTryon (Guo et al., 2025)). The position of the images (i.e., left/right) was randomized to prevent bias. Participants were asked to select the image that best represented a high-fidelity, in-shop version of the garment worn by the model, considering texture preservation, structural integrity, and overall realism.
Results. We collected a total of 1,920 pairwise judgments. The results, summarized in Table 8, demonstrate a strong preference for our method. TEMU-VTOFF outperforms MGT with a win rate of 75.77% and Any2AnyTryon with a win rate of 77.74%. These results strongly corroborate our quantitative analyses (particularly the DISTS and FID scores), confirming that TEMU-VTOFF produces results that are perceptually superior to state-of-the-art methods.
Table 8: Pairwise comparison showing the percentage of times participants preferred TEMU-VTOFF over the competing method. “Not Sure” indicates cases where participants found the quality of generated images indistinguishable.
Comparison Ours Wins (%) Not Sure (%) Competitor Wins (%)
TEMU-VTOFF vs. MGT 75.77 7.85 16.38
TEMU-VTOFF vs. Any2AnyTryon 77.74 5.64 16.62
Figure 8: Additional qualitative results of TEMU-VTOFF and competitors on VITON-HD (Choi et al., 2021), with per-sample metrics. DISTS emphasizes structural differences between images better than LPIPS, confirming its higher correlation with human judgments.
E Discussion and Limitations
@@
Our method demonstrates strong performance and generalization, yet it inherits some inner problems of foundational models such as Stable Diffusion 3 (Esser et al., 2024). Although we improve the rendering of large logos and text, the model still struggles with fine-grained details, including complex texture patterns, small printed text, and the correct reproduction of small objects such as buttons. Moreover, as mentioned in the main paper, reconstruction is less reliable for lower-body garments than for upper-body items or dresses, likely due to class imbalance in the Dress Code dataset. For completeness, we show a set of failure cases in Fig. 13 and Fig. 14, on sample images from Dress Code (Cui et al., 2021) and VITON-HD (Lee et al., 2022), respectively.
We further analyze how the adopted perceptual metrics correlate with the presence of visual artifacts in the generated images (Fig. 8). While quantitative comparisons are reported as averages over the full test set, inspecting per-sample metric values is particularly informative in VTOFF, as it always lives in a paired setting. Edge cases, such as missing garment components or incorrect structural details, are often critical, and VTOFF naturally provides paired person-garment samples. In this context, LPIPS and DISTS play an important role, as both measure image-to-image distances. It is therefore essential to verify that these metrics respond reliably to detail discrepancies and appropriately penalize weaker baselines.
For each example in Fig. 8, we display three images: the ground-truth target, the output of TEMU-VTOFF, and the output of TryOffAnyone (Xarchakos and Koukopoulos, 2024). We present four representative comparisons. The first two (columns 1-3) show cases where TEMU-VTOFF preserves fine garment details that are lost by the competitor. The remaining two (columns 4-6) contrast accurate samples from our method with misaligned or rotated outputs produced by TryOffAnyone. For each pair, we report the corresponding per-sample DISTS and LPIPS values, along with the percentage improvement of our results over those of TryOffAnyone. In situations where a clear qualitative gap exists, DISTS consistently reflects the expected difference, whereas LPIPS often fails to penalize severe distortions and occasionally even assigns worse scores to the better-performing method (e.g., row 2, column 5). These observations provide empirical evidence supporting our choice to include DISTS as part of the overall evaluation protocol. This observation is consistent with findings from DreamSim (Fu et al., 2023), as discussed in Sec. A.2.
F LLM Usage
In this work, we employ LLMs (specifically Qwen2.5-VL) to extract garment-related textual descriptions, which serve as conditioning signals for generation. Beyond this, LLMs were employed solely for minor language refinement. They did not contribute to the design of experiments, the analysis of results, or the generation of scientific content.
Figure 9: Additional qualitative results of TEMU-VTOFF and competitors on Dress Code (Morelli et al., 2022).
Figure 10: Additional qualitative results of TEMU-VTOFF and competitors on VITON-HD (Choi et al., 2021).
Figure 11: Additional qualitative results showing the contribution of each component in TEMU-VTOFF on Dress Code (Morelli et al., 2022) images.
Figure 12: Inputs used to generate the target garment with TEMU-VTOFF, using sample images from Dress Code (Cui et al., 2021)